Copper Mountain Group​ (U.S.).The Copper Mountain​ Group, a private equity firm headquartered in​ Boulder, Colorado, borrows £4,500,000 for one year at 7.625% interest.

a.What is the dollar cost of this debt if the pound depreciates from $2.0270​/£ to $1.9450​/£ over the​ year?

b.What is the dollar cost of this debt if the pound appreciates from $2.0270​/£ to $2.1610​/£ over the​ year?
Answer #1

(a) Original Borrowing =  £4,500,000 , Initial Exchange Rate = $ 2.027 / £, $ Value of Borrowing = 2.027 x 4500000 = $ 9121500

Interest Rate = 7.625 %

£ Value of Borrowing after 1 year = 4500000 x 1.07625 = £ 4843125

Exchange Rate after 1 year = $ 1.945 / £

$ Value of Borrowing after 1 year = 4843125 x 1.945 = $ 9419878

Let the $ cost of borrowing ($ interest rate) be R

Therefore, 9121500 x (1+R) = 9419878

R = [9419878 / 9121500] - 1 = 0.032712 or 3.27 %

(b) Original Borrowing =  £4,500,000 , Initial Exchange Rate = $ 2.027 / £, $ Value of Borrowing = 2.027 x 4500000 = $ 9121500

Interest Rate = 7.625 %

£ Value of Borrowing after 1 year = 4500000 x 1.07625 = £ 4843125

Exchange Rate after 1 year = $ 2.161 / £

$ Value of Borrowing after 1 year = 4843125 x 2.161 = $ 10465993.13  

Let the $ cost of borrowing ($ interest rate) be R

Therefore, 9121500 x (1+R) = 9419878

R = [10465993.13 / 9121500] - 1 = 0.14739 or 14.74 %